# Explicit unit distance graphs with exponential chromatic number and arbitrary girth

### v1

@@ Abstract
We prove that there exist unit distance graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.107+o(1))^{d}$ that have arbitrarily large girth. This improves upon a series of results due to Kupavskii; Sagdeev; and Sagdeev and Raigorodskii and gives the first bound in which the base of the exponent does not tend to one with the girth. In addition, our construction can be made explicit which allows us to answer in a strong form a question of Kupavskii.
Our arguments show graphs of large chromatic number and high girth exist in a number of other geometric settings including diameter graphs and orthogonality graphs.
@@ 1 Introduction
Chromatic number and girth are two of the perhaps most fundamental graph parameters, the study of which has attracted a vast amount of attention over the years. The chromatic number of a graph is the minimum number of colours needed to colour its vertices so that no two adjacent vertices receive the same colour (such a colouring is referred to as proper), and the girth of a graph is equal to the length of a shortest cycle in the graph. One of the most classical directions of study of these two parameters explores the relation between them. In particular, the question of whether one can obtain some amount of control of the chromatic number in terms of the girth has a very long and illustrious history. Indeed, at first glance, it is highly non-trivial to construct graphs with very high chromatic number, where this can not be witnessed locally. The local behaviour is captured quite well by the assumption of having large girth since this is equivalent to saying that the graph needs to locally look like a forest and any forest has chromatic number at most two.
@@
The study of the chromatic number of a variety of geometric graphs is equally as classical a topic. For example, the Hadwiger-Nelson problem [19], dating back to 1945, asks for the maximum chromatic number of a unit distance graph in $\mathbb{R}^{d}$ while the classical Borsuk Problem [3], dating back to 1933, asks for the maximum chromatic number of a diameter graph in $\mathbb{R}^{d}$. Here a graph is said to be a unit distance graph in $\mathbb{R}^{d}$ if we can assign its vertices to distinct points in $\mathbb{R}^{d}$ in such a way that any pair of adjacent vertices are assigned to points at a unit distance. A graph is said to be a diameter graph in $\mathbb{R}^{d}$ if we can assign its vertices to distinct points in $\mathbb{R}^{d}$ in such a way that any pair of adjacent vertices are assigned to a pair of points at the maximum distance. We note that in both settings one may assume graphs involved are finite via the classical de-Bruijn-Erdős Compactness Theorem [6].
In the unit distance setting, in their remarkable 1981 paper essentially introducing the algebraic method in combinatorics, Frankl and Wilson [17] proved a conjecture of Erdős that there exist unit distance graphs in $\mathbb{R}^{d}$ with chromatic number exponential in $d$, improving upon a sequence of progressively better polynomial bounds (see e.g. the survey [44] for the full history). The Frankl-Wilson construction has been slightly improved since by Raigorodskii who holds the current record of $\chi(G)\geq(1.239+o(1))^{d}$ for a unit distance graph in $\mathbb{R}^{d}$.
In the diameter graph setting no bound beyond $\chi(G)=d+1$ obtained by the regular simplex was known. It was widely believed and is referred to as the Borsuk Conjecture that $\chi(G)\leq d+1$ for any diameter graph in $\mathbb{R}^{d}$, based in part on numerous results establishing this for special configurations of points. This was disproved in a stunning fashion by Kahn and Kalai in [21] who constructed diameter graphs with chromatic number growing exponentially in $\sqrt{d}$. There have been several improvements in the base, as well as a considerable effort in trying to prove any non-trivial bounds for small $d$, but constructing a diameter graph with exponential chromatic number remains a very interesting open problem. The current record is $\chi(G)\geq(1.225+o(1))^{\sqrt{d}}$ for $G$ a diameter graph in $\mathbb{R}^{d}$, due to Raigorodskii [42]. See the survey [43] for a more complete history and plenty of related problems.
@@
Kupavskii [23] in 2012 proved that for each positive integer $k$ there exists some $c_{k}>1$ such that there are unit distance graphs in $\mathbb{R}^{d}$ with girth at least $k$ and chromatic number at least $(c_{k}+o(1))^{d}$. Alon and Kupavskii [2] reiterate the question of what can be said about the chromatic number of such graphs. Sagdeev [47] gave the first explicit estimate for $c_{k}$, namely that $c_{k}\geq 1+1/{k^{4+o(1)}}$. The dependency of $c_{k}$ on $k$ was subsequently improved to $c_{k}\geq 1+1/{k^{2+o(1)}}$ in [48] and by a further constant factor (implicit in $1/{k^{2+o(1)}}$) by Sagdeev and Raigorodskii in [46]. We prove a first universal exponential lower bound, without any dependency on $k$.
@@
There exist unit distance graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.107+o(1))^{d}$ that have arbitrarily large girth.
Our bound even improves the best-known bounds for small values of $k$, e.g. in perhaps the most classical case of lower bounding the chromatic number of unit distance triangle-free graphs in $\mathbb{R}^{d}$ the previously best bound was $(1.058+o(1))^{d}$ [7]. The bounds on chromatic number of unit distance graphs in which we forbid a clique of size $k$ rather than all short cycles has also been considered, see [7, 25]. The bound in Theorem 1 improves the best-known lower bounds for the chromatic number of unit distance graphs in $\mathbb{R}^{d}$ with clique number at most $7$.
Quite a bit of focus over the years was afforded to finding explicit constructions of high girth and high chromatic number graphs, ever since the first such construction was found by Lovász in 1966 [26], in large part motivated by the idea that such graphs should exhibit somewhat random-like behaviour so could be useful in a variety of derandomisation arguments. Explicit constructions of graphs with large girth and chromatic number are now known in some geometric settings, including as box and line intersection graphs in $\mathbb{R}^{3}$ [4], as disjointedness graphs of short polygonal chains in the plane [35], and as tangency graphs of circles in the plane [5]. Our construction behind Theorem 1 is actually more flexible than previous, weaker ones in the unit distance graph setting and can be made explicit. This answers in quite a strong form a question of Kupavskii who asked to prove explicitly his result that unit distance graphs in $\mathbb{R}^{d}$ with girth at least $k$ and chromatic number at least $(c_{k}+o(1))^{d}$ exist for some $c_{k}>1$.
@@
Another setting, underlying and connecting the previous two results, is that of orthogonality graphs. Here a graph is said to be an orthogonality graph in $\mathbb{R}^{d}$ if we can assign its vertices to distinct vectors in $\mathbb{R}^{d}$ in such a way that any pair of adjacent vertices are assigned to orthogonal vectors.
@@
There exist orthogonality graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.107+o(1))^{d}$ that have arbitrarily large girth.
@@ 2 A self-contained argument
The edge set will be obtained by randomly subsampling orthogonal pairs of vectors in $V$. Let us however first introduce some notation. We let $V^{\prime}:=\{\textbf{v}\mid\textbf{v}\in\{\pm 1\}^{d/2},\textbf{v}_{d/2}=1,\sum_{i=1}^{d/2}\textbf{v}_{i}=0\}$ and for $\textbf{v}\in V^{\prime},j\in[m]$ we let $\textbf{v}(j)=\left(\textbf{v}\cdot\cos\frac{2\pi j}{m},\textbf{v}\cdot\sin\frac{2\pi j}{m}\right)$. So that $V=\{\textbf{v}(j)\mid\textbf{v}\in V^{\prime},j\in[m]\}$. One can think of $V$ as consisting of $m$ spaced out rotations of each $\textbf{v}\in V^{\prime}$ in the plane defined by $(\textbf{v},\textbf{0})$ and $(\textbf{0},\textbf{v})$ in $\mathbb{R}^{d}$.
The key property of $V$ is that if v is orthogonal to u for $\textbf{v},\textbf{u}\in V^{\prime}$ then $\textbf{v}(i)$ is orthogonal to $\textbf{u}(j)$ for any $i,j\in[m]$. So if one denotes by $G^{\prime}$ the graph with vertex set $V^{\prime}$ and edge set being orthogonal pairs in $V^{\prime}$, then the blow-up graph $G^{\prime m}$ (which is the graph obtained from $G^{\prime}$ by replacing each vertex with an independent set of size $m$ and each edge with a $K_{m,m}$ subgraph between the two corresponding independent sets of size $m$) is an orthogonality graph in $\mathbb{R}^{d}$ with vertex set $V$. Armed with this observation our plan is to show that $\chi(G^{\prime})$ is exponential in $d$ and that thanks to that if we subsample the edges of $G^{\prime m}$ we can preserve the same chromatic number while removing all cycles of length shorter than $g$.
@@
    $\frac{|V^{\prime}|}{2\binom{4p}{p-1}}=\frac{\binom{4p}{2p}}{4\binom{4p}{p-1}}\geq\left({27}/{16}+o(1)\right)^{p}\geq\left(1.067+o(1)\right)^{d}$
    colours to properly colour $G^{\prime}$, as claimed. ∎
@@
We now turn to the second part of the argument, namely increasing the girth. Let $k=\chi(G^{\prime})$ and let $m=2^{8pg}\cdot(4k)^{4g}\cdot(k-1)^{2}$. As we already argued $G^{\prime m}$ is an orthogonality graph in $\mathbb{R}^{d}$. Let $n=m\cdot 2^{4p}\geq|V(G^{\prime m})|.$ Let $G_{0}\sim G^{\prime m}(q)$ be the random graph obtained by including every edge of $G^{\prime m}$ in $G_{0}$ with probability $q:=1/n^{1-1/(2g)}$, independently from all other edges. The expected number of cycles of length $\ell$ in $G_{0}$ is at most $\binom{n}{\ell}\cdot\frac{\ell!}{2\ell}\cdot q^{\ell}\leq(nq)^{\ell}/2=n^{\ell/(2g)}/2$. So, the expected number of cycles of length at most $g$ is at most $\sum_{\ell=3}^{g}n^{\ell/(2g)}/2\leq n^{1/2}.$ By Markov’s inequality, we can conclude that the probability there are more than $2\sqrt{n}$ cycles in $G_{0}$ is less than $1/2$.
@@
On the other hand, if we write $m^{\prime}:=\frac{m}{k-1}$ we want to show that with probability less than $1/2$ there exist adjacent $v,u\in V(G^{\prime})$ and a pair of subsets $V,U$ of $m^{\prime}$ vertices of $G^{\prime m}$, corresponding to $v$ and $u$ respectively with at most $2\sqrt{n}$ edges between $V$ and $U$. To see that note that there are at most $e(G^{\prime})\cdot\binom{m}{m^{\prime}}^{2}$ choices for $V$ and $U$ and the chance there are at most $2\sqrt{n}$ edges between them is at most $\binom{m^{\prime 2}}{2\sqrt{n}}(1-q)^{m^{\prime 2}-2\sqrt{n}}$. So the desired statement holds by a union bound as
$e(G^{\prime})\cdot\binom{m}{m^{\prime}}^{2}\cdot\binom{m^{\prime 2}}{2\sqrt{n}}(1-q)^{m^{\prime 2}-2\sqrt{n}}\leq 2^{8p}\cdot 2^{2m}\cdot 2^{4\sqrt{n}\log m^{\prime}}\cdot 2^{-qm^{\prime 2}+1}<\frac{1}{2},$
where the final inequality follows since $qm^{\prime 2}\geq m^{1+1/(2g)}/(2^{4p}k^{2})\geq 4m\geq 8p+2m+4\sqrt{n}\log m+2.$
@@
So there exists an outcome in which graph $G_{0}$ has at most $2\sqrt{n}$ cycles of length up to $g$ and for any adjacent $v,u\in V(G^{\prime})$ and for any pair of subsets $V,U$ of $m^{\prime}$ vertices of $G^{\prime m}$, corresponding to $v$ and $u$ respectively, there are more than $2\sqrt{n}$ edges between $V$ and $U$. We obtain our graph $G$ by removing an arbitrary edge from every cycle of length up to $g$ in $G_{0}$ to obtain a graph with girth at least $g$ and with the property that there is an edge between any pair of subsets $V,U$ of $m^{\prime}$ vertices of $G^{\prime m}$, corresponding to $v$ and $u$ respectively.
@@
Suppose that $\chi(G)<\chi(G^{\prime})=k$. This means there exists a colouring $c:V(G)\to[k-1]$ such that there are no adjacent vertices with the same colour. Now colour every vertex $v$ of $G^{\prime}$ by the majority colour according to $c$ appearing on the $m$ vertices corresponding to $v$ in $V(G^{\prime m})=V(G).$ This produces a $(k-1)$-colouring of $G^{\prime}$ which can not be proper, as $\chi(G^{\prime})=k$, and hence there must exist adjacent $v,u\in V(G^{\prime})$ given the same colour. By definition of the colouring, there exist two subsets $V,U$ of $\frac{m}{k-1}=m^{\prime}$ vertices of $G^{\prime m}$, corresponding to $v$ and $u$ respectively, all given the same colour in $c$. But there exists an edge between these $V$ and $U$ by our construction, which contradicts the definition of $c$.
@@
This completes the proof of Theorem 3 with a somewhat weaker base of the exponential (being 1.067). Next, we obtain Theorem 1 and Theorem 2 with the same somewhat weaker base of the exponentials. The same arguments also show that Theorem 3 implies Theorem 1 and Theorem 2.
@@
Since all the vertices in $V$ have the same norm $||\textbf{v}(j)||^{2}=||\textbf{v}||^{2}\left(\cos^{2}\frac{2\pi j}{m}+\sin^{2}\frac{2\pi j}{m}\right)=||\textbf{v}||^{2}=4p$ we know that any orthogonal pair $\textbf{v}(i),\textbf{v}(j)$ satisfy $||\textbf{v}(i)-\textbf{v}(j)||^{2}=8p$. So after scaling all the vectors by a factor of $\sqrt{8p}$ we obtain our graph $G$ is also a unit distance graph, proving Theorem 1 with a smaller base of the exponential.

### v3

# Geometric graphs with exponential chromatic number and arbitrary girth

@@ Abstract
We prove that there exist unit distance graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.074+o(1))^{d}$ that have arbitrarily large girth. This improves upon a series of results due to Kupavskii; Sagdeev; and Sagdeev and Raigorodskii and gives the first bound in which the base of the exponent does not tend to one with the girth. In addition, our construction can be made explicit which allows us to answer in a strong form a question of Kupavskii.
Our arguments show graphs of large chromatic number and high girth exist in a number of other geometric settings, including diameter graphs and orthogonality graphs.
@@ 1 Introduction
Chromatic number and girth are perhaps two of the most fundamental graph parameters, the study of which has attracted a vast amount of attention over the years. The chromatic number of a graph $G$ (denoted by $\chi(G)$) is the minimum number of colours needed to colour its vertices so that no two adjacent vertices receive the same colour (such a colouring is referred to as proper), and the girth of a graph is equal to the length of a shortest cycle in the graph. One of the most classical directions of study of these two parameters explores the relation between them. In particular, the question of whether one can obtain some amount of control of the chromatic number in terms of the girth has a very long and illustrious history. Indeed, at first glance, it is highly non-trivial to construct graphs with very high chromatic number, where this can not be witnessed locally. The local behaviour is captured quite well by the assumption of having large girth since this is equivalent to saying that the graph needs to locally look like a forest and any forest has chromatic number at most two.
@@
The study of the chromatic number of a variety of geometric graphs is equally as classical a topic. For example, the Hadwiger-Nelson problem [19], dating back to 1944, asks for the maximum chromatic number of a unit distance graph in $\mathbb{R}^{d}$ while the classical Borsuk Problem [3], dating back to 1933, asks for the maximum chromatic number of a diameter graph in $\mathbb{R}^{d}$. Here a graph is said to be a unit distance graph in $\mathbb{R}^{d}$ if we can assign its vertices to distinct points in $\mathbb{R}^{d}$ in such a way that any pair of adjacent vertices are assigned to points at a unit distance [^1] We note that here and throughout the paper we do not insist that this assignment is faithful, namely we do not require that any pair of points at distance one actually correspond to an edge in the graph, we discuss this distinction further in Section 4.. A graph is said to be a diameter graph in $\mathbb{R}^{d}$ if we can assign its vertices to distinct points in $\mathbb{R}^{d}$ in such a way that any pair of adjacent vertices are assigned to a pair of points at the maximum distance. We note that in both settings one may assume graphs involved are finite via the classical de-Bruijn-Erdős Compactness Theorem [7].
In the unit distance setting, in their remarkable 1981 paper essentially introducing the algebraic method in combinatorics, Frankl and Wilson [17] proved a conjecture of Erdős that there exist unit distance graphs in $\mathbb{R}^{d}$ with chromatic number exponential in $d$, improving upon a sequence of progressively better polynomial bounds (see e.g. the survey [42] for the full history). The Frankl-Wilson construction has been slightly improved since by Raigorodskii [41] who holds the current record of $\chi(G)\geq(1.239+o(1))^{d}$ for a unit distance graph in $\mathbb{R}^{d}$. The best known upper bound $(3+o(1))^{d}$ is due to Larman and Rogers [26] (see also [40] for a more recent simpler proof due to Prosanov).
In the diameter graph setting, no bound beyond $\chi(G)=d+1$, obtained by the regular simplex, was known. It was widely believed and is referred to as the Borsuk Conjecture that $\chi(G)\leq d+1$ for any diameter graph in $\mathbb{R}^{d}$, based in part on numerous results establishing this for special configurations of points. This was disproved in a stunning fashion by Kahn and Kalai in [21] who constructed diameter graphs with chromatic number growing exponentially in $\sqrt{d}$. There have been several improvements in the base, as well as a considerable effort in trying to prove any non-trivial bounds for small $d$, but constructing a diameter graph with exponential chromatic number remains a very interesting open problem. The current record is $\chi(G)\geq(1.225+o(1))^{\sqrt{d}}$ for $G$ a diameter graph in $\mathbb{R}^{d}$, due to Raigorodskii [46]. The best known upper bound here is $(3/2)^{(1+o(1))d/2}$ due to Schramm [52], see also [4] for a different proof due to Bourgain and Lindenstrauss. See the survey [47] for a more complete history and plenty of related problems. For example, in [23] the authors construct counterexamples to Borsuk Conjecture in any large enough dimension which are in a certain sense very different compared to the Kahn-Kalai construction.
@@
Kupavskii [24] in 2012 proved that for each positive integer $g$ there exists some $c_{g}>1$ such that there are unit distance graphs in $\mathbb{R}^{d}$ with girth at least $g$ and chromatic number at least $(c_{g}+o(1))^{d}$. Alon and Kupavskii [2] reiterate the question of what can be said about the chromatic number of such graphs. Sagdeev [49] gave the first explicit estimate for $c_{g}$, namely that $c_{g}\geq 1+1/{g^{4+o(1)}}$. The dependency of $c_{g}$ on $g$ was subsequently improved to $c_{g}\geq 1+1/{g^{2+o(1)}}$ in [50] and by a further constant factor (implicit in $1/{g^{2+o(1)}}$) by Sagdeev and Raigorodskii in [51]. We prove a first universal exponential lower bound, without any dependency on $g$.
@@
There exist unit distance graphs in $\mathbb{R}^{d}$ with chromatic number at least $(1.074+o(1))^{d}$ that have arbitrarily large girth.
Our bound even improves the best-known bounds for small values of $g$, e.g. in perhaps the most classical case of lower bounding the chromatic number of unit distance triangle-free graphs in $\mathbb{R}^{d}$ the previously best bound was $(1.058+o(1))^{d}$ [8].
Quite a bit of focus over the years was afforded to finding explicit constructions of high girth and high chromatic number graphs, ever since the first such construction was found by Lovász in 1966 [27], in large part motivated by the idea that such graphs should exhibit somewhat random-like behaviour so could be useful in a variety of derandomisation arguments. Explicit constructions of graphs with large girth and chromatic number are now known in some geometric settings, including as box and line intersection graphs in $\mathbb{R}^{3}$ [5], as disjointedness graphs of short polygonal chains in the plane [36], and as tangency graphs of circles in the plane [6]. Our construction behind Theorem 1 is actually more flexible than previous, weaker ones in the unit distance graph setting and can be made explicit. This answers in quite a strong form a question of Kupavskii who asked to prove explicitly his result that unit distance graphs in $\mathbb{R}^{d}$ with girth at least $g$ and chromatic number at least $(c_{g}+o(1))^{d}$ exist for some $c_{g}>1$.
@@
Another setting, underlying and connecting the previous two results, is that of orthogonality graphs. Here a graph is said to be an orthogonality graph in $S\subseteq\mathbb{R}^{d}$ if we can assign its vertices to distinct vectors in $S$ in such a way that any pair of adjacent vertices are assigned to orthogonal vectors. In this paper, we will mostly work with orthogonality graphs over spheres, here we denote by $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$ the $d-1$ dimensional unit sphere consisting of all vectors of unit length in $\mathbb{R}^{d}$.
@@
There exist orthogonality graphs in $\mathbb{S}^{d-1}$ with chromatic number at least $(1.074+o(1))^{d}$ that have arbitrarily large girth.
@@ 2 A self-contained argument
The edge set will be obtained by randomly subsampling orthogonal pairs of vectors in $V$. Let us however first introduce some notation. We let $V^{\prime}:=\{\textbf{v}\mid\textbf{v}\in\{\pm 1\}^{d/2},\textbf{v}_{d/2}=1,\sum_{i=1}^{d/2}\textbf{v}_{i}=0\}$ and for $\textbf{v}\in V^{\prime},j\in[m]$ we let $\textbf{v}(j)=\left(\textbf{v}\cdot\cos\frac{2\pi j}{m},\textbf{v}\cdot\sin\frac{2\pi j}{m}\right)$. So that $V=\{\textbf{v}(j)\mid\textbf{v}\in V^{\prime},j\in[m]\}\subseteq\frac{d}{2}\cdot\mathbb{S}^{d-1}$. One can think of $V$ as consisting of $m$ spaced out rotations of each $\textbf{v}\in V^{\prime}$ in the plane defined by $(\textbf{v},\textbf{0})$ and $(\textbf{0},\textbf{v})$ in $\mathbb{R}^{d}$.
The key property of $V$ is that if v is orthogonal to u for $\textbf{v},\textbf{u}\in V^{\prime}$ then $\textbf{v}(i)$ is orthogonal to $\textbf{u}(j)$ for any $i,j\in[m]$. So if one denotes by $G^{\prime}$ the graph with vertex set $V^{\prime}$ and edge set being orthogonal pairs in $V^{\prime}$, then the blow-up graph $G^{\prime m}$ (which is the graph obtained from $G^{\prime}$ by replacing each vertex with an independent set of size $m$ and each edge with a $K_{m,m}$ subgraph between the two corresponding independent sets of size $m$) is an orthogonality graph in $\frac{d}{2}\cdot\mathbb{S}^{d-1}$ with vertex set $V$ (so also in $\mathbb{S}^{d-1}$ after scaling by a factor of $2/d$). Armed with this observation our plan is to show that $\chi(G^{\prime})$ is exponential in $d$ and that thanks to that if we subsample the edges of $G^{\prime m}$ we can preserve the same chromatic number while removing all cycles of length shorter than $g$.
@@
    Let $S\subseteq V^{\prime}$ contain no orthogonal pair of vectors. We first show that this implies $|S|\leq 2\binom{4p}{p-1}$. To see this let us define polynomials $f_{\textbf{v}}:\{\pm 1\}^{4p}\to\mathbb{F}_{p}$ for each $\textbf{v}\in V^{\prime}$ as
    $f_{\textbf{v}}(\textbf{x}):=\prod_{i=1}^{p-1}(\langle\textbf{x},\textbf{v}\rangle-i)\bmod p.$
@@
    $\frac{|V^{\prime}|}{2\binom{4p}{p-1}}=\frac{\binom{4p}{2p}}{4\binom{4p}{p-1}}=\left({27}/{16}+o(1)\right)^{p}\geq\left(1.067+o(1)\right)^{d}$
    colours to properly colour $G^{\prime}$, as claimed. We note that the term $27/16$ arises from Stirling’s approximation which we used to estimate the ratio of the two binomials. ∎
@@
We now turn to the second part of the argument, namely increasing the girth. Let $k=\chi(G^{\prime})$ and let $m=2^{8pg}\cdot(4k)^{4g}\cdot(k-1)^{2}$. As we already argued $G^{\prime m}$ is an orthogonality graph in $\mathbb{S}^{d-1}$. Let $n=m\cdot 2^{4p}\geq|V(G^{\prime m})|.$ Let $G_{0}\sim G^{\prime m}(q)$ be the random graph obtained by including every edge of $G^{\prime m}$ in $G_{0}$ with probability $q:=1/n^{1-1/(2g)}$, independently from all other edges. The expected number of cycles of length $\ell$ in $G_{0}$ is at most $\binom{n}{\ell}\cdot\frac{\ell!}{2\ell}\cdot q^{\ell}\leq(nq)^{\ell}/2=n^{\ell/(2g)}/2$. So, the expected number of cycles of length at most $g$ is at most $\sum_{\ell=3}^{g}n^{\ell/(2g)}/2\leq n^{1/2}$ (using that $n^{1/(2g)}\geq 2$ which holds for our choice of $n$). By Markov’s inequality, we can conclude that the probability there are more than $2\sqrt{n}$ cycles in $G_{0}$ is less than $1/2$.
@@
On the other hand, if we write $m^{\prime}:=\frac{m}{k-1}$ we want to show that with probability less than $1/2$ there exist adjacent $\textbf{v},\textbf{u}\in V(G^{\prime})$ and a pair of subsets $W,U$ of $m^{\prime}$ vertices of $G^{\prime m}$, corresponding to v and u respectively with at most $2\sqrt{n}$ edges between $W$ and $U$. To see that note that there are at most $e(G^{\prime})\cdot\binom{m}{m^{\prime}}^{2}$ choices for $W$ and $U$ and the chance there are at most $2\sqrt{n}$ edges between them is at most $\binom{m^{\prime 2}}{2\sqrt{n}}(1-q)^{m^{\prime 2}-2\sqrt{n}}$. So the desired statement holds by a union bound as
$e(G^{\prime})\cdot\binom{m}{m^{\prime}}^{2}\cdot\binom{m^{\prime 2}}{2\sqrt{n}}(1-q)^{m^{\prime 2}-2\sqrt{n}}\leq 2^{8p}\cdot 2^{2m}\cdot 2^{4\sqrt{n}\log_{2}m^{\prime}}\cdot 2^{-qm^{\prime 2}+1}<\frac{1}{2},$
where the final inequality follows since $qm^{\prime 2}\geq m^{1+1/(2g)}/(2^{4p}k^{2})>4m\geq 8p+2m+4\sqrt{n}\log_{2}m+2.$
@@
So there exists an outcome in which graph $G_{0}$ has at most $2\sqrt{n}$ cycles of length up to $g$ and for any adjacent $\textbf{v},\textbf{u}\in V(G^{\prime})$ and for any pair of subsets $W,U$ of $m^{\prime}$ vertices of $G^{\prime m}$, corresponding to v and u respectively, there are more than $2\sqrt{n}$ edges between $W$ and $U$. We obtain our graph $G$ by removing an arbitrary edge from every cycle of length up to $g$ in $G_{0}$ to obtain a graph with girth at least $g$ and with the property that there is an edge between any pair of subsets $W,U$ of $m^{\prime}$ vertices of $G^{\prime m}$, corresponding to v and u respectively.
@@
Suppose that $\chi(G)<\chi(G^{\prime})=k$. This means there exists a colouring $c:V(G)\to[k-1]$ such that there are no adjacent vertices with the same colour. Now colour every vertex v of $G^{\prime}$ by the majority colour according to $c$ appearing on the $m$ vertices corresponding to v in $V(G^{\prime m})=V(G).$ This produces a $(k-1)$-colouring of $G^{\prime}$ which can not be proper, as $\chi(G^{\prime})=k$, and hence there must exist adjacent $\textbf{v},\textbf{u}\in V(G^{\prime})$ given the same colour. By definition of the colouring, there exist two subsets $W,U$ of $\frac{m}{k-1}=m^{\prime}$ vertices of $G^{\prime m}$, corresponding to v and u respectively, all given the same colour in $c$. But there exists an edge between these $W$ and $U$ by our construction, which contradicts the definition of $c$.
@@
This completes the proof of Theorem 3 with a somewhat weaker base of the exponential (being 1.067). Next, we obtain Theorem 1 and Theorem 2 with weaker base of the exponentials. The same arguments also show that Theorem 3 implies Theorem 1 and Theorem 2 with bounds as stated.
We start with the classical, folklore observation that any orthogonality graph on a sphere $\mathbb{S}^{d-1}$ is also a unit distance graph in $\mathbb{R}^{d}$. Note that all the vertices in our set $W$ have the same norm
$||\textbf{v}(j)||^{2}=||\textbf{v}||^{2}\left(\cos^{2}\frac{2\pi j}{m}+\sin^{2}\frac{2\pi j}{m}\right)=||\textbf{v}||^{2}=4p.$
This implies that any orthogonal pair of vectors $\textbf{v}(i),\textbf{v}(j)$ in $V$ satisfy
$||\textbf{v}(i)-\textbf{v}(j)||^{2}=||\textbf{v}(i)||^{2}+||\textbf{v}(j)||^{2}-2\langle\textbf{v}(i),\textbf{v}(j)\rangle=8p.$ (1)
@@
So after scaling all the vectors by a factor of $\sqrt{8p}$ we obtain our graph $G$ is also a unit distance graph, proving Theorem 1 with a smaller base of the exponential. We note that, in general, if we are given an orthogonality graph in $\mathbb{R}^{d}$ we can scale all of its vectors to have the same norm without impacting the orthogonality relations making it an orthogonality graph on $\mathbb{S}^{d-1}$ so long as no vectors scale into the same point. This is the reason we insist on working over $\mathbb{S}^{d-1}$ rather than $\mathbb{R}^{d}$.
[…]
$\langle\textbf{v}\otimes\textbf{v},\textbf{w}\otimes\textbf{w}\rangle=\sum_{i,j}\textbf{v}_{i}\textbf{v}_{j}\textbf{w}_{i}\textbf{w}_{j}=\sum_{i}\textbf{v}_{i}\textbf{w}_{i}\cdot\sum_{j}\textbf{v}_{j}\textbf{w}_{j}=\langle\textbf{v},\textbf{w}\rangle^{2}\geq 0.$